Write the slope-intercept equation of the line l that goes through (12, 3) and has a slopeof 1/3
.

Found 2 solutions by Alan3354, checkley77:
Answer by Alan3354(69443)

Write the slope-intercept equation of the line l that goes through (12, 3) and has a slope of 1/3
y - y1 = m(x - x1)
y - 3 = 1/3*(x - 12)
3y - 9 = x - 12
x - 3y = 3
The y intercept is where x = 0
So, 0 - 3y = 3
y = -1 is the y-intercept
So y = x/3 - 1 (same as the equation above)

Answer by checkley77(12844)

slope=1/3 through (12,3)
Y=mX+b WHERE m=SLOPE & b=Y INTERCEPT.
3=12*1/3+b
3=4+b
b=3-4
b=-1 THE Y INTERCEPT.
THUS THE EQUATION IS:
Y=X/3-1